OpenStudy (anonymous):

can someone integrate 1/4x^2

OpenStudy (lgbasallote):

@Rogue answered this already :O

OpenStudy (anonymous):

I asked this question earlier, but made a mistake and wrote differentiate, instead of integrate.

OpenStudy (lgbasallote):

he wrote the integral

OpenStudy (anonymous):

How do i go back to my first problem? I didn't see it.

OpenStudy (lgbasallote):

go to your profile (you can bress that bar beside the f button) then questions asked

OpenStudy (anonymous):

ok thank you. can you tell me if my answer is correct? I got -1/16x^2 when finding the integration of 1/(2x)^3

OpenStudy (anonymous):

@lgbasallote , how to you "bress" something?

OpenStudy (anonymous):

*how do

OpenStudy (anonymous):

@lgbasallote. Can you tell me if my answer is correct? I got -1/16x^2 when finding the integration of 1/(2x)^3.

OpenStudy (lgbasallote):

let me think -_-

OpenStudy (lgbasallote):

i'll check by differentiating -1/16x^2 should come out as 1/(2x)^3 \[\frac{d}{dx}(\frac{-x^{-2}}{16}) = \frac{2x^{-3}}{16} = \frac{2}{16x^3} = \frac{1}{8x^3} = \frac{1}{(2x)^3} \checkmark\]

OpenStudy (anonymous):

Thank you.